OpenStudy (anonymous):

Solve the given equation. 3e^x + 1 = 5 a. 0.2877 b. 2 c. 2.877 d. -0.2877

OpenStudy (nincompoop):

do you know what the e stands for or its value?

OpenStudy (shamim):

3e^x=4

OpenStudy (nurali):

3e^x + 1 = 5 3e^x=5-1 3e^x=4 e^x=4/3 e^x=1.33 Taking ln Both Side lne^x=ln(1.33) xlne=0.123 lne=1 x=0.2877

OpenStudy (anonymous):

e=2.71 its 0.2877

OpenStudy (anonymous):

x=4/3e^ x=4/14.1267 now x=0.2877

OpenStudy (anonymous):

thank you @rathanreddy & @Nurali

OpenStudy (nurali):

Anytime.

OpenStudy (anonymous):

e=2.71 its 0.2877.

OpenStudy (nincompoop):

e≠2.71 you still haven't answered what e stands for and its value, @elsa1995 I don't think you're learning anything by just going by how people are solving it.

OpenStudy (nincompoop):

someone is hungry for a medal :D